# Well-posedness of a generalized Stokes operator on domains with cylindrical ends via layer-potentials

### 11 May 2026 v1

@@ 7.2. Green formulas and energy estimates for indicial families
In view of further applications to manifolds with cylindrical ends, we want to extend the Green and representation formulas and the energy estimates of Section 2 (especially, Subsection 2.3) to the case of the indicial family $\widehat{\boldsymbol{\Xi}}$ of $\boldsymbol{\Xi}:=\boldsymbol{\Xi}_{V,V_{0}}$ (see Definition 5.26). Thus, in this subsection, we continue to consider the cylinder $M=\mathfrak{S}\times\mathbb{R}$, where $\mathfrak{S}$ is a closed manifold. The results of this section will be used for $\mathfrak{S}$ not necessarily connected, so we will not assume that $\mathfrak{S}$ is connected, although the general case can immediately be reduced to the connected case. Also, we let $\color[rgb]{1,0,1}\widehat{V}(x^{\prime})={\mathcal{R}_{\infty}}V(x^{\prime},t)=V(x^{\prime},t)\geq 0$ and $\color[rgb]{1,0,1}\widehat{V}_{0}(x^{\prime})={\mathcal{R}_{\infty}}V_{0}(x^{\prime},t)=V_{0}(x^{\prime},t)\geq 0$ be smooth, translation invariant functions (this notation is justified by Equation (5.6)). Again, this choice of $V$ and $V_{0}$ satisfies all our previous assumptions. (Note that in some of our previous papers, the notation $\tilde{V}$ and $\tilde{V}_{0}$ was used instead of $\color[rgb]{1,0,1}\widehat{V}$ and $\color[rgb]{1,0,1}\widehat{V}_{0}$. However, in this paper, the notation $\tilde{u}$ means $\tilde{u}(x):=u(-x)$.)
@@
Also, in this subsection, we let $\Omega^{\prime}\subset\mathfrak{S}$ be a smooth open subset with boundary $\Gamma^{\prime}:=\partial\Omega^{\prime}$ (a smooth manifold). Our results are formulated in general, but there is no loss of generality to assume that $\Omega^{\prime}$ is connected (that is, a smooth domain). Also, recall that, in this section, $\color[rgb]{1,0,1}\widehat{V}(x^{\prime})=V(x^{\prime},t)$ and $\color[rgb]{1,0,1}\widehat{V}_{0}(x^{\prime})=V_{0}(x^{\prime},t)$, as motivated by Equation (5.6). In particular, $V=\color[rgb]{1,0,1}\widehat{V}$ and $V_{0}=\color[rgb]{1,0,1}\widehat{V}_{0}$ are translation invariant. The main equation of Notation 2.10 becomes.
(7.12) $\begin{gathered}U\mathrel{\,:=\,}(\boldsymbol{u}\ \ p)^{\top}\,,\ W\mathrel{\,:=\,}(\boldsymbol{w}\ \ q)^{\top}\in H^{1}(\Omega^{\prime};TM)\oplus L^{2}(\Omega^{\prime})\,,\\ \qquad\mathfrak{v}\mathrel{\,:=\,}({\color[rgb]{1,0,1}\widehat{V}}\boldsymbol{u},\boldsymbol{w})_{\Omega^{\prime}}-({\color[rgb]{1,0,1}\widehat{V}_{0}}p,q)_{\Omega^{\prime}}\,,\quad\mbox{and}\\ B_{\Omega^{\prime},\tau}(U,W)\mathrel{\,:=\,}2(\widehat{\operatorname{Def}}\,(\tau)\boldsymbol{u},\widehat{\operatorname{Def}}\,(\tau)\boldsymbol{w})_{\Omega^{\prime}}+(\widehat{\nabla}^{*}(\tau)\boldsymbol{u},q)_{\Omega^{\prime}}+(p,\widehat{\nabla}^{*}(\tau)\boldsymbol{w})_{\Omega^{\prime}}+\mathfrak{v}\,.\end{gathered}$
@@
(7.13) $\widehat{\boldsymbol{\Xi}}(\tau)\,=\,\left(\begin{array}[]{cc}2\widehat{\operatorname{Def}}^{*}\,(\tau)\widehat{\operatorname{Def}}\,(\tau)+\color[rgb]{1,0,1}\widehat{V}&\widehat{\nabla}(\tau)\\ \widehat{\nabla}^{*}(\tau)&-\color[rgb]{1,0,1}\widehat{V}_{0}\end{array}\right)\,$,
where, we recall, we are identifying $\color[rgb]{1,0,1}\widehat{V}=\widehat{V}(\tau)$ and $\color[rgb]{1,0,1}\widehat{V}_{0}=\widehat{V}_{0}(\tau)$. (The formula (7.13) is valid for any manifold with straight cylindrical ends, not just for our manifold $M:=\mathfrak{S}\times\mathbb{R}$.) The last proposition gives then the following lemma analogous to Corollary 2.14. Recall that, we are assuming that $V,V_{0}\geq 0$. Also, recall that $\Omega^{\prime}\subset\mathfrak{S}$ is a smooth domain.
@@
Let $\tau\in\mathbb{R}$, let $\color[rgb]{1,0,1}\widehat{V}\in\mathcal{C}^{\infty}(\mathfrak{S}\times\mathbb{R};\operatorname{End}(TM))^{\mathbb{R}}\simeq\mathcal{C}^{\infty}(\mathfrak{S};TM)$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\in\mathcal{C}^{\infty}(\mathfrak{S}\times\mathbb{R})^{\mathbb{R}}\simeq\mathcal{C}^{\infty}(\mathfrak{S})$ be non-negative (i.e., $\color[rgb]{1,0,1}\widehat{V},\color[rgb]{1,0,1}\widehat{V}_{0}\geq 0$), and let
@@
$\widehat{\operatorname{Def}}\,(\tau)\boldsymbol{u}\,=\,0\,,\quad\color[rgb]{1,0,1}\widehat{V}\boldsymbol{u}\,=\,0\,,\quad\widehat{\nabla}^{*}(\tau)\boldsymbol{u}\,=\,0\,,\quad\widehat{\nabla}(\tau)p\,=\,0\,,\quad\mbox{and}\quad\color[rgb]{1,0,1}\widehat{V}_{0}p\,=\,0\,.$
@@
of Proposition 7.5. Let $\operatorname{Re}(z)$ denote the real part of $z\in\mathbb{C}$. Then $\operatorname{Re}\big[-(\widehat{\nabla}^{*}(\tau)\boldsymbol{u},p)_{\Omega^{\prime}}+(p,\widehat{\nabla}^{*}(\tau)\boldsymbol{u})_{\Omega^{\prime}}\big]=0$. Let $\mathfrak{v}:=(\color[rgb]{1,0,1}\widehat{V}\boldsymbol{u},\boldsymbol{u})_{\Omega^{\prime}}+(\color[rgb]{1,0,1}\widehat{V}_{0}p,p)_{\Omega^{\prime}}$, as before. Together with the definition of $B_{\Omega^{\prime},\tau}$ in Equation (7.12), this gives
@@
= $\,=\,2(\widehat{\operatorname{Def}}\,(\tau)\boldsymbol{u},\widehat{\operatorname{Def}}\,(\tau)\boldsymbol{u})_{\Omega^{\prime}}+(\color[rgb]{1,0,1}\widehat{V}\boldsymbol{u},\boldsymbol{u})_{\Omega^{\prime}}+(\color[rgb]{1,0,1}\widehat{V}_{0}p,p)_{\Omega^{\prime}}\,.$
All three terms in the last sum are non-negative, so each of them equals zero. Therefore $\widehat{\operatorname{Def}}\,(\tau)\boldsymbol{u}=0$, $\color[rgb]{1,0,1}\widehat{V}\boldsymbol{u}=0$, and $\color[rgb]{1,0,1}\widehat{V}_{0}p=0$. We also have
$0\,=\,\widehat{\boldsymbol{\Xi}}(\tau)U\,=\,\left(\begin{array}[]{c}{2\widehat{\operatorname{Def}}^{*}\,(\tau)\widehat{\operatorname{Def}}\,(\tau)\boldsymbol{u}+\color[rgb]{1,0,1}\widehat{V}\boldsymbol{u}+\widehat{\nabla}(\tau)p}\\ {\widehat{\nabla}^{*}(\tau)\boldsymbol{u}-\color[rgb]{1,0,1}\widehat{V}_{0}p}\end{array}\right)\,=\,\left(\begin{array}[]{c}{\widehat{\nabla}(\tau)p}\\ {\widehat{\nabla}^{*}(\tau)u}\end{array}\right)\,.$
@@
Let $\color[rgb]{1,0,1}\widehat{V},\color[rgb]{1,0,1}\widehat{V}_{0}\geq 0$ and let $U=(\boldsymbol{u}\ \ p)^{\top}\in H^{2}(\Omega^{\prime};TM)\oplus H^{1}(\Omega^{\prime})$ satisfy $\widehat{\boldsymbol{\Xi}}(0)U=0$ in $\Omega^{\prime}$ and $(\widehat{\boldsymbol{T}}_{\boldsymbol{\nu}}(0)U,\boldsymbol{u})_{\Gamma^{\prime}}=0$ (as in Lemma 7.6, but for $\tau=0$).
@@
    If $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\Omega^{\prime}$ (Definition 2.13), then $p=0$ on $\Omega^{\prime}$.
@@
      $\color[rgb]{1,0,1}\widehat{V}\succ 0$ on $\Omega^{\prime}$,
@@
Lemma 7.6 gives $\widehat{\nabla}(0)p=0$ in $\Omega^{\prime}$. The assumption that $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ in $\Omega$ together with Proposition 7.3 (stating the $L^{2}$-unique continuation property of $\widehat{\nabla}(0)$ in $\Omega^{\prime}$) then give $p=0$. Similarly, Lemma 7.6 gives $\widehat{\operatorname{Def}}\,(0)\boldsymbol{u}=0$. Our assumptions allow us to use Proposition 7.4 (the $L^{2}$-unique continuation property of $\widehat{\operatorname{Def}}\,(0)$) to conclude that $\boldsymbol{u}=0$. This completes the proof. ∎
@@
Let $\color[rgb]{1,0,1}\widehat{V},\color[rgb]{1,0,1}\widehat{V}_{0}\geq 0$ and let $U=(\boldsymbol{u}\ \ p)^{\top}\in H^{2}(\mathfrak{S};TM)\oplus H^{1}(\mathfrak{S})$ satisfy $\widehat{\boldsymbol{\Xi}}(0)U=0$ in $\mathfrak{S}$.
@@
    If $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\mathfrak{S}$, then $p=0$ on $\mathfrak{S}$.
@@
    Similarly, if $\color[rgb]{1,0,1}\widehat{V}\succ 0$ on $\mathfrak{S}$, then $\boldsymbol{u}=0$ in $\mathfrak{S}$.
@@
Let $\color[rgb]{1,0,1}\widehat{V},\color[rgb]{1,0,1}\widehat{V}_{0}\geq 0$, $\tau\neq 0$, and $U^{\top}=(\boldsymbol{u}\ \ p)\in H^{2}(\Omega^{\prime};TM)\oplus H^{1}(\Omega^{\prime})$ satisfy $\widehat{\boldsymbol{\Xi}}(\tau)U=0$ in $\Omega^{\prime}$ and $(\widehat{\boldsymbol{T}}_{\boldsymbol{\nu}}(\tau)U,\boldsymbol{u})^{\prime}=0$ (the same assumptions as in Lemma 7.6, except that $\tau\neq 0$ in this corollary). Then $U=0$.
@@
Let $\color[rgb]{1,0,1}\widehat{V},\color[rgb]{1,0,1}\widehat{V}_{0}\geq 0$, $\tau\neq 0$, and $U^{\top}=(\boldsymbol{u}\ \ p)\in H^{2}(\mathfrak{S};TM)\oplus H^{1}(\mathfrak{S})$ satisfy $\widehat{\boldsymbol{\Xi}}(\tau)U=0$. Then $U=0$.
@@ 7.3. The invertibility of $\boldsymbol{\Xi}$ on a cylinder
Recall from Theorem 6.15 that the Fredholm property of an operator depends on the invertibility of its limit operator, which acts on the cylinder associated to $M$. The limit operator of $\boldsymbol{\Xi}_{V,V_{0}}$ is $\boldsymbol{\Xi}_{\color[rgb]{1,0,1}\widehat{V},\color[rgb]{1,0,1}\widehat{V}_{0}}$. It is the purpose of this subsection to establish the invertibility of $\boldsymbol{\Xi}_{\color[rgb]{1,0,1}\widehat{V},\color[rgb]{1,0,1}\widehat{V}_{0}}$ acting on the associated cylinder under the assumption that $\color[rgb]{1,0,1}\widehat{V}\succ 0$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\mathfrak{S}$.
We thus continue to assume, in this subsection, that $M$ is the cylinder $M=\mathfrak{S}\times\mathbb{R}$ with $\mathfrak{S}$ a closed manifold and with a product metric. We also assume that $V(x^{\prime},t)=\color[rgb]{1,0,1}\widehat{V}(x^{\prime})$ and that $V_{0}(x^{\prime},t)=\color[rgb]{1,0,1}\widehat{V}_{0}(x^{\prime})$, and hence $V$ and $V_{0}$ are thus translation invariant. We also assume that they are both $\succ 0$ on $M$ (see Definition 2.13). This condition is equivalent to $\color[rgb]{1,0,1}\widehat{V}\succ 0$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\mathfrak{S}$.
@@
Let us assume that $M=\mathfrak{S}\times\mathbb{R}$ is a cylinder and that $\color[rgb]{1,0,1}\widehat{V}\succ 0$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\mathfrak{S}$. Let $\boldsymbol{s}=\boldsymbol{t}=(1,0)$. Then $\widehat{\boldsymbol{\Xi}}(0):=\widehat{\boldsymbol{\Xi}}_{V,V_{0}}(0)$ is invertible and $\widehat{\boldsymbol{\Xi}}(0)^{-1}\in\Psi^{[-\boldsymbol{s}-\boldsymbol{t}]}(\mathfrak{S};TM\oplus\mathbb{C})$.
@@
The second Green formula of Proposition 7.5 shows that $\widehat{\boldsymbol{\Xi}}(0)$ is self-adjoint, and hence it has index zero. Therefore, in order to show that $\widehat{\boldsymbol{\Xi}}(0)$ induces a bijection between the above spaces, it suffices to show that it is injective. Let then $U=\left(\begin{array}[]{c}{\boldsymbol{u}}\\ {p}\end{array}\right)=(\,\boldsymbol{u}\ \ p\,)^{\top}\in L^{2}(\mathfrak{S};TM\oplus\mathbb{C})$ be such that $\widehat{\boldsymbol{\Xi}}(0)U=0$. The ellipticity of $\widehat{\boldsymbol{\Xi}}(0)$ guarantees then that $U\in H^{\infty}(\mathfrak{S};TM\oplus\mathbb{C})$, by elliptic regularity (see, for example, Theorem 15.3.30 in [63]). Assumption 8.3 allow us to use Corollary 7.8 to conclude that $U=0$. Therefore $\widehat{\boldsymbol{\Xi}}(0)$ is invertible as an operator between the indicated spaces. The property that its inverse belongs to the space $\Psi^{[-\boldsymbol{s}-\boldsymbol{t}]}(\mathfrak{S};TM\oplus\mathbb{C})$ follows from the work of Beals and others (see Theorem 7.4 in [67] for more details). ∎
The case $\tau\neq 0$ is completely similar, but only requires the weaker assumptions $\color[rgb]{1,0,1}\widehat{V}:={\mathcal{R}_{\infty}}(V)\geq 0$ and $\color[rgb]{1,0,1}\widehat{V}_{0}:={\mathcal{R}_{\infty}}(V_{0})\geq 0$.
@@
Let us assume that $M=\mathfrak{S}\times\mathbb{R}$ is a cylinder, that $\tau\neq 0$, and that $\color[rgb]{1,0,1}\widehat{V}_{0},\color[rgb]{1,0,1}\widehat{V}\geq 0$. Let $\boldsymbol{s}=\boldsymbol{t}=(1,0)$. Then $\widehat{\boldsymbol{\Xi}}(\tau)$ is invertible and $\widehat{\boldsymbol{\Xi}}(\tau)^{-1}\in\Psi^{[-\boldsymbol{s}-\boldsymbol{t}]}(\mathfrak{S};TM\oplus\mathbb{C})$.
@@
The proof is almost the same as that of Proposition 7.11 (but easier). The only difference is that, instead of Corollary 7.8, we use Corollary 7.10, which only requires that $\color[rgb]{1,0,1}\widehat{V}$ and $\color[rgb]{1,0,1}\widehat{V}_{0}$ be non-negative. ∎
@@
Let us assume that $M=\mathfrak{S}\times\mathbb{R}$ with $\mathfrak{S}$ compact and that $\color[rgb]{1,0,1}\widehat{V}$ and $\color[rgb]{1,0,1}\widehat{V}_{0}$ are non-negative. Then, for any $\tau,r\in\mathbb{R}$, $\tau\neq 0$, we have an isomorphism
@@
Let us assume that $M=\mathfrak{S}\times\mathbb{R}$, that $V$ and $V_{0}$ are translation invariant and that $\color[rgb]{1,0,1}\widehat{V}\succ 0$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\mathfrak{S}$. Then the operator

### 28 May 2026 v2

@@ 7.2. Green formulas and energy estimates for indicial families
In view of further applications to manifolds with cylindrical ends, we want to extend the Green and representation formulas and the energy estimates of Section 2 (especially, Subsection 2.3) to the case of the indicial family $\widehat{\boldsymbol{\Xi}}$ of $\boldsymbol{\Xi}:=\boldsymbol{\Xi}_{V,V_{0}}$ (see Definition 5.26). Thus, in this subsection, we continue to consider the cylinder $M=\mathfrak{S}\times\mathbb{R}$, where $\mathfrak{S}$ is a closed manifold. The results of this section will be used for $\mathfrak{S}$ not necessarily connected, so we will not assume that $\mathfrak{S}$ is connected, although the general case can immediately be reduced to the connected case. Also, we let $\widehat{V}(x^{\prime})={\mathcal{R}_{\infty}}V(x^{\prime},t)=V(x^{\prime},t)\geq 0$ and $\widehat{V}_{0}(x^{\prime})={\mathcal{R}_{\infty}}V_{0}(x^{\prime},t)=V_{0}(x^{\prime},t)\geq 0$ be smooth, translation invariant functions (this notation is justified by Equation (5.6)). Again, this choice of $V$ and $V_{0}$ satisfies all our previous assumptions. (Note that in some of our previous papers, the notation $\tilde{V}$ and $\tilde{V}_{0}$ was used instead of $\widehat{V}$ and $\widehat{V}_{0}$. However, in this paper, the notation $\tilde{u}$ means $\tilde{u}(x):=u(-x)$.)
@@
Also, in this subsection, we let $\Omega^{\prime}\subset\mathfrak{S}$ be a smooth open subset with boundary $\Gamma^{\prime}:=\partial\Omega^{\prime}$ (a smooth manifold). Our results are formulated in general, but there is no loss of generality to assume that $\Omega^{\prime}$ is connected (that is, a smooth domain). Also, recall that, in this section, $\widehat{V}(x^{\prime})=V(x^{\prime},t)$ and $\widehat{V}_{0}(x^{\prime})=V_{0}(x^{\prime},t)$, as motivated by Equation (5.6). In particular, $V=\widehat{V}$ and $V_{0}=\widehat{V}_{0}$ are translation invariant. The main equation of Notation 2.10 becomes.
(7.12) $\begin{gathered}U\mathrel{\,:=\,}(\boldsymbol{u}\ \ p)^{\top}\,,\ W\mathrel{\,:=\,}(\boldsymbol{w}\ \ q)^{\top}\in H^{1}(\Omega^{\prime};TM)\oplus L^{2}(\Omega^{\prime})\,,\\ \qquad\mathfrak{v}\mathrel{\,:=\,}({\widehat{V}}\boldsymbol{u},\boldsymbol{w})_{\Omega^{\prime}}-({\widehat{V}_{0}}p,q)_{\Omega^{\prime}}\,,\quad\mbox{and}\\ B_{\Omega^{\prime},\tau}(U,W)\mathrel{\,:=\,}2(\widehat{\operatorname{Def}}\,(\tau)\boldsymbol{u},\widehat{\operatorname{Def}}\,(\tau)\boldsymbol{w})_{\Omega^{\prime}}+(\widehat{\nabla}^{*}(\tau)\boldsymbol{u},q)_{\Omega^{\prime}}+(p,\widehat{\nabla}^{*}(\tau)\boldsymbol{w})_{\Omega^{\prime}}+\mathfrak{v}\,.\end{gathered}$
@@
(7.13) $\widehat{\boldsymbol{\Xi}}(\tau)\,=\,\left(\begin{array}[]{cc}2\widehat{\operatorname{Def}}^{*}\,(\tau)\widehat{\operatorname{Def}}\,(\tau)+\widehat{V}&\widehat{\nabla}(\tau)\\ \widehat{\nabla}^{*}(\tau)&-\widehat{V}_{0}\end{array}\right)\,$,
where, we recall, we are identifying $\widehat{V}=\widehat{V}(\tau)$ and $\widehat{V}_{0}=\widehat{V}_{0}(\tau)$. (The formula (7.13) is valid for any manifold with straight cylindrical ends, not just for our manifold $M:=\mathfrak{S}\times\mathbb{R}$.) The last proposition gives then the following lemma analogous to Corollary 2.14. Recall that, we are assuming that $V,V_{0}\geq 0$. Also, recall that $\Omega^{\prime}\subset\mathfrak{S}$ is a smooth domain.
@@
Let $\tau\in\mathbb{R}$, let $\widehat{V}\in\mathcal{C}^{\infty}(\mathfrak{S}\times\mathbb{R};\operatorname{End}(TM))^{\mathbb{R}}\simeq\mathcal{C}^{\infty}(\mathfrak{S};TM)$ and $\widehat{V}_{0}\in\mathcal{C}^{\infty}(\mathfrak{S}\times\mathbb{R})^{\mathbb{R}}\simeq\mathcal{C}^{\infty}(\mathfrak{S})$ be non-negative (i.e., $\widehat{V},\widehat{V}_{0}\geq 0$), and let
@@
$\widehat{\operatorname{Def}}\,(\tau)\boldsymbol{u}\,=\,0\,,\quad\widehat{V}\boldsymbol{u}\,=\,0\,,\quad\widehat{\nabla}^{*}(\tau)\boldsymbol{u}\,=\,0\,,\quad\widehat{\nabla}(\tau)p\,=\,0\,,\quad\mbox{and}\quad\widehat{V}_{0}p\,=\,0\,.$
@@
of Proposition 7.5. Let $\operatorname{Re}(z)$ denote the real part of $z\in\mathbb{C}$. Then $\operatorname{Re}\big[-(\widehat{\nabla}^{*}(\tau)\boldsymbol{u},p)_{\Omega^{\prime}}+(p,\widehat{\nabla}^{*}(\tau)\boldsymbol{u})_{\Omega^{\prime}}\big]=0$. Let $\mathfrak{v}:=(\widehat{V}\boldsymbol{u},\boldsymbol{u})_{\Omega^{\prime}}+(\widehat{V}_{0}p,p)_{\Omega^{\prime}}$, as before. Together with the definition of $B_{\Omega^{\prime},\tau}$ in Equation (7.12), this gives
@@
= $\,=\,2(\widehat{\operatorname{Def}}\,(\tau)\boldsymbol{u},\widehat{\operatorname{Def}}\,(\tau)\boldsymbol{u})_{\Omega^{\prime}}+(\widehat{V}\boldsymbol{u},\boldsymbol{u})_{\Omega^{\prime}}+(\widehat{V}_{0}p,p)_{\Omega^{\prime}}\,.$
All three terms in the last sum are non-negative, so each of them equals zero. Therefore $\widehat{\operatorname{Def}}\,(\tau)\boldsymbol{u}=0$, $\widehat{V}\boldsymbol{u}=0$, and $\widehat{V}_{0}p=0$. We also have
$0\,=\,\widehat{\boldsymbol{\Xi}}(\tau)U\,=\,\left(\begin{array}[]{c}{2\widehat{\operatorname{Def}}^{*}\,(\tau)\widehat{\operatorname{Def}}\,(\tau)\boldsymbol{u}+\widehat{V}\boldsymbol{u}+\widehat{\nabla}(\tau)p}\\ {\widehat{\nabla}^{*}(\tau)\boldsymbol{u}-\widehat{V}_{0}p}\end{array}\right)\,=\,\left(\begin{array}[]{c}{\widehat{\nabla}(\tau)p}\\ {\widehat{\nabla}^{*}(\tau)u}\end{array}\right)\,.$
@@
Let $\widehat{V},\widehat{V}_{0}\geq 0$ and let $U=(\boldsymbol{u}\ \ p)^{\top}\in H^{2}(\Omega^{\prime};TM)\oplus H^{1}(\Omega^{\prime})$ satisfy $\widehat{\boldsymbol{\Xi}}(0)U=0$ in $\Omega^{\prime}$ and $(\widehat{\boldsymbol{T}}_{\boldsymbol{\nu}}(0)U,\boldsymbol{u})_{\Gamma^{\prime}}=0$ (as in Lemma 7.6, but for $\tau=0$).
@@
    If $\widehat{V}_{0}\succ 0$ on $\Omega^{\prime}$ (Definition 2.13), then $p=0$ on $\Omega^{\prime}$.
@@
      $\widehat{V}\succ 0$ on $\Omega^{\prime}$,
@@
Lemma 7.6 gives $\widehat{\nabla}(0)p=0$ in $\Omega^{\prime}$. The assumption that $\widehat{V}_{0}\succ 0$ in $\Omega$ together with Proposition 7.3 (stating the $L^{2}$-unique continuation property of $\widehat{\nabla}(0)$ in $\Omega^{\prime}$) then give $p=0$. Similarly, Lemma 7.6 gives $\widehat{\operatorname{Def}}\,(0)\boldsymbol{u}=0$. Our assumptions allow us to use Proposition 7.4 (the $L^{2}$-unique continuation property of $\widehat{\operatorname{Def}}\,(0)$) to conclude that $\boldsymbol{u}=0$. This completes the proof. ∎
@@
Let $\widehat{V},\widehat{V}_{0}\geq 0$ and let $U=(\boldsymbol{u}\ \ p)^{\top}\in H^{2}(\mathfrak{S};TM)\oplus H^{1}(\mathfrak{S})$ satisfy $\widehat{\boldsymbol{\Xi}}(0)U=0$ in $\mathfrak{S}$.
@@
    If $\widehat{V}_{0}\succ 0$ on $\mathfrak{S}$, then $p=0$ on $\mathfrak{S}$.
@@
    Similarly, if $\widehat{V}\succ 0$ on $\mathfrak{S}$, then $\boldsymbol{u}=0$ in $\mathfrak{S}$.
@@
Let $\widehat{V},\widehat{V}_{0}\geq 0$, $\tau\neq 0$, and $U^{\top}=(\boldsymbol{u}\ \ p)\in H^{2}(\Omega^{\prime};TM)\oplus H^{1}(\Omega^{\prime})$ satisfy $\widehat{\boldsymbol{\Xi}}(\tau)U=0$ in $\Omega^{\prime}$ and $(\widehat{\boldsymbol{T}}_{\boldsymbol{\nu}}(\tau)U,\boldsymbol{u})^{\prime}=0$ (the same assumptions as in Lemma 7.6, except that $\tau\neq 0$ in this corollary). Then $U=0$.
@@
Let $\widehat{V},\widehat{V}_{0}\geq 0$, $\tau\neq 0$, and $U^{\top}=(\boldsymbol{u}\ \ p)\in H^{2}(\mathfrak{S};TM)\oplus H^{1}(\mathfrak{S})$ satisfy $\widehat{\boldsymbol{\Xi}}(\tau)U=0$. Then $U=0$.
@@ 7.3. The invertibility of $\boldsymbol{\Xi}$ on a cylinder
Recall from Theorem 6.15 that the Fredholm property of an operator depends on the invertibility of its limit operator, which acts on the cylinder associated to $M$. The limit operator of $\boldsymbol{\Xi}_{V,V_{0}}$ is $\boldsymbol{\Xi}_{\widehat{V},\widehat{V}_{0}}$. It is the purpose of this subsection to establish the invertibility of $\boldsymbol{\Xi}_{\widehat{V},\widehat{V}_{0}}$ acting on the associated cylinder under the assumption that $\widehat{V}\succ 0$ and $\widehat{V}_{0}\succ 0$ and $\widehat{V}_{0}\succ 0$ on $\mathfrak{S}$.
We thus continue to assume, in this subsection, that $M$ is the cylinder $M=\mathfrak{S}\times\mathbb{R}$ with $\mathfrak{S}$ a closed manifold and with a product metric. We also assume that $V(x^{\prime},t)=\widehat{V}(x^{\prime})$ and that $V_{0}(x^{\prime},t)=\widehat{V}_{0}(x^{\prime})$, and hence $V$ and $V_{0}$ are thus translation invariant. We also assume that they are both $\succ 0$ on $M$ (see Definition 2.13). This condition is equivalent to $\widehat{V}\succ 0$ and $\widehat{V}_{0}\succ 0$ on $\mathfrak{S}$.
@@
Let us assume that $M=\mathfrak{S}\times\mathbb{R}$ is a cylinder and that $\widehat{V}\succ 0$ and $\widehat{V}_{0}\succ 0$ on $\mathfrak{S}$. Let $\boldsymbol{s}=\boldsymbol{t}=(1,0)$. Then $\widehat{\boldsymbol{\Xi}}(0):=\widehat{\boldsymbol{\Xi}}_{V,V_{0}}(0)$ is invertible and $\widehat{\boldsymbol{\Xi}}(0)^{-1}\in\Psi^{[-\boldsymbol{s}-\boldsymbol{t}]}(\mathfrak{S};TM\oplus\mathbb{C})$.
@@
The second Green formula of Proposition 7.5 shows that $\widehat{\boldsymbol{\Xi}}(0)$ is self-adjoint, and hence it has index zero. Therefore, in order to show that $\widehat{\boldsymbol{\Xi}}(0)$ induces a bijection between the above spaces, it suffices to show that it is injective. Let then $U=\left(\begin{array}[]{c}{\boldsymbol{u}}\\ {p}\end{array}\right)=(\,\boldsymbol{u}\ \ p\,)^{\top}\in L^{2}(\mathfrak{S};TM\oplus\mathbb{C})$ be such that $\widehat{\boldsymbol{\Xi}}(0)U=0$. The ellipticity of $\widehat{\boldsymbol{\Xi}}(0)$ guarantees then that $U\in H^{\infty}(\mathfrak{S};TM\oplus\mathbb{C})$, by elliptic regularity (see Theorem 6.16 or [83, Theorem 15.3.30] or [87, Proposition 6.5]). Assumption 8.3 allow us to use Corollary 7.8 to conclude that $U=0$. Therefore $\widehat{\boldsymbol{\Xi}}(0)$ is invertible as an operator between the indicated spaces. The property that its inverse belongs to the space $\Psi^{[-\boldsymbol{s}-\boldsymbol{t}]}(\mathfrak{S};TM\oplus\mathbb{C})$ follows from the work of Beals and others (see Theorem 7.4 in [87] for more details). ∎
The case $\tau\neq 0$ is completely similar, but only requires the weaker assumptions $\widehat{V}:={\mathcal{R}_{\infty}}(V)\geq 0$ and $\widehat{V}_{0}:={\mathcal{R}_{\infty}}(V_{0})\geq 0$.
@@
Let us assume that $M=\mathfrak{S}\times\mathbb{R}$ is a cylinder, that $\tau\neq 0$, and that $\widehat{V}_{0},\widehat{V}\geq 0$. Let $\boldsymbol{s}=\boldsymbol{t}=(1,0)$. Then $\widehat{\boldsymbol{\Xi}}(\tau)$ is invertible and $\widehat{\boldsymbol{\Xi}}(\tau)^{-1}\in\Psi^{[-\boldsymbol{s}-\boldsymbol{t}]}(\mathfrak{S};TM\oplus\mathbb{C})$.
@@
The proof is almost the same as that of Proposition 7.11 (but easier). The only difference is that, instead of Corollary 7.8, we use Corollary 7.10, which only requires that $\widehat{V}$ and $\widehat{V}_{0}$ be non-negative. ∎
@@
Let us assume that $M=\mathfrak{S}\times\mathbb{R}$ with $\mathfrak{S}$ compact and that $\widehat{V}$ and $\widehat{V}_{0}$ are non-negative. Then, for any $\tau,r\in\mathbb{R}$, $\tau\neq 0$, we have an isomorphism
@@
Let us assume that $M=\mathfrak{S}\times\mathbb{R}$, that $V$ and $V_{0}$ are translation invariant and that $\widehat{V}\succ 0$ and $\widehat{V}_{0}\succ 0$ on $\mathfrak{S}$. Then the operator